please factor the following and explain, thank you
x^2-25 over x+5
and
4y^2+16y+16

Answer by alka001(14)

x^2-25/x+5
=(x-5)(x+5)/x+5..........formula x^2-25=x^2-5^2=(x-5)(x+5)
=x-5

4y^2+16y+16
sum of numbers=16;product of numbers=16*4=64
factors of 64 whose sum is 16 are8,8
4y^2+8y+8y+16
=4y(y+2)+8(y+2)
=(4y+8)(y+2)